OpenStudy (anonymous):

Someone may answer this question. abs(x)+x=0

OpenStudy (jiteshmeghwal9):

abs=absolute value ?

OpenStudy (jiteshmeghwal9):

@viniterranova ?

OpenStudy (anonymous):

Yes

OpenStudy (jiteshmeghwal9):

|x|=x whether it is positive or negative so, x+x=0 2x=0 x=0/2=0

OpenStudy (jiteshmeghwal9):

gt it @viniterranova ??

OpenStudy (anonymous):

Yes. Thanks

OpenStudy (jiteshmeghwal9):

yw :)